OpenStudy (anonymous):

Please tell me if f(x) = x^3 + 0.04x^2 + 3 is odd, even or niether (then tell me how you found out)

OpenStudy (anonymous):

Hey jim!!

jimthompson5910 (jim_thompson5910):

hey

jimthompson5910 (jim_thompson5910):

Hint: if f(-x) = f(x), then the function is even

OpenStudy (anonymous):

I'm doin' good in class. Your help on my math packet got me some good brownie points :)

jimthompson5910 (jim_thompson5910):

if f(-x) = -f(x), then the function is odd

OpenStudy (anonymous):

right. But I'm not sure how to show it

jimthompson5910 (jim_thompson5910):

so first, let's find f(-x)

OpenStudy (anonymous):

does f(-x) = -x^3 -0.04x + 3

jimthompson5910 (jim_thompson5910):

f(x) = x^3 + 0.04x^2 + 3 f(-x) = (-x)^3 + 0.04(-x)^2 + 3 f(-x) = -x^3 + 0.04x^2 + 3

jimthompson5910 (jim_thompson5910):

keep in mind that (-x)^2 = x^2

jimthompson5910 (jim_thompson5910):

so does f(-x) = f(x)???

OpenStudy (anonymous):

but not (-x)^3

jimthompson5910 (jim_thompson5910):

(-x)^3 = (-x)*(-x)*(-x) = -x^3

OpenStudy (anonymous):

so f(-x) does not = f(x)

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

which means that f is not an even function

OpenStudy (anonymous):

so it's neither ?

jimthompson5910 (jim_thompson5910):

now what about f(-x) = -f(x)?

OpenStudy (anonymous):

F IS NOT EVEN A FUNCTION? Omg dude

OpenStudy (anonymous):

lol

jimthompson5910 (jim_thompson5910):

huh

OpenStudy (anonymous):

just kidding around

OpenStudy (anonymous):

f(-x) = -x^3 - 0.04x - 3?

jimthompson5910 (jim_thompson5910):

we can refer to f(x) as f, I've seen it plenty of times

OpenStudy (anonymous):

I'm not sure how to show f(-x) = -f(x)

jimthompson5910 (jim_thompson5910):

negate everything in f(x) to get -f(x)

jimthompson5910 (jim_thompson5910):

-f(x) = -1*f(x)

jimthompson5910 (jim_thompson5910):

So if f(x) = x^3 + 0.04x^2 + 3, then -f(x) = -1*f(x) = -1(x^3 + 0.04x^2 + 3)

jimthompson5910 (jim_thompson5910):

then you distribute and simplify

OpenStudy (anonymous):

neither instantly and with your eyeballs

OpenStudy (anonymous):

so it would be -x^3+0.04x^2+3 = -x^3-0.04x^2-3. so it is neither.

OpenStudy (anonymous):

\[f(x) = x^3 + 0.04x^2 + 3 \] has both even and odd powers, so neither

jimthompson5910 (jim_thompson5910):

yep you got it

OpenStudy (anonymous):

thanks jim :)

jimthompson5910 (jim_thompson5910):

yw